OpenStudy (ksaimouli):

suppose a ball is thrown vertically upward.Eight seconds later it returns to its point of release.What is initial velocity of the ball.

OpenStudy (anonymous):

dont know ??? :D

OpenStudy (ksaimouli):

dont know

OpenStudy (ksaimouli):

what formula to use

OpenStudy (anonymous):

1 min !

OpenStudy (anonymous):

V=s/t

OpenStudy (ksaimouli):

i think u r wrong @RaoAbdulAhad

OpenStudy (ajprincess):

Use \(v=u+at\) v=-u -u=u-gt gt=2u u=gt/2 =10*8/2

OpenStudy (ksaimouli):

why v=-u

OpenStudy (anonymous):

??

OpenStudy (ajprincess):

when it reaches the ground it has the same velocity as the initial velocity. Since the direction is opposite u put - in front of it

OpenStudy (ksaimouli):

ok thx

OpenStudy (ajprincess):

yw.:)